# On optimal solutions of classical and sliced Wasserstein GANs with non-Gaussian data

### v1

@@ IV-A Proof for Theorem 1
As (33) is equivalent to (35), Theorem 1 directly shows that $\Theta^{*}=(\theta_{1}^{*},\theta_{2}^{*})\in\mathbb{R}\times\mathbb{R}_{+}$ as in (9) is the unique minimizer of (33). Thanks to (31) and (9),
@@
$=\int_{\mathbb{R}}(x-\mathbb{E}[X])^{2}d\mu(x)$
$\hskip 14.45377pt-2\theta^{*}_{2}\int_{\mathbb{R}}(x-\mathbb{E}[X])\Big{(}\Psi^{-1}(F_{\mu}(x))-\mathbb{E}[h(Z)]\Big{)}d\mu(x)$
$\hskip 50.58878pt+(\theta^{*}_{2})^{2}\int_{\mathbb{R}}\Big{(}\Psi^{-1}(F_{\mu}(x))-\mathbb{E}[h(Z)]\Big{)}^{2}d\mu(x)$
where the fourth equality follows from that $\Psi^{-1}(F_{\mu}(X))$ and $h(Z)$ have identical distribution and the last equality is due to the definition of $\theta^{*}_{2}$ in (9). ∎
@@
Suppose that $\mu$ has no atom, the cdf $\Psi$ of $h(Z)$ is continuous and strictly increasing, and $\operatorname{Var}(h(Z))>0$. The sub-problem
$\min_{\theta_{1}\in\mathbb{R},\theta_{2}\leq 0}\int_{\mathbb{R}}|x-t^{\Theta}(x)|^{2}d\mu(x)$ (44)
has a unique minimizer $\Theta^{*}=(\theta_{1}^{*},\theta_{2}^{*})\in\mathbb{R}\times\mathbb{R}_{+}$, defined by
@@
Proof.
$\begin{split}\eta^{*}&=\frac{\operatorname{Cov}\left(X,\Psi^{-1}_{0}(F_{\mu}(X))\right)}{\operatorname{Var}(-h(Z))}=-\frac{\operatorname{Cov}\left(X,\Psi^{-1}(1-F_{\mu}(X))\right)}{\operatorname{Var}(h(Z))}\\ \theta^{*}_{1}&=\mathbb{E}[X]-\eta^{*}\mathbb{E}\left[-h(Z)\right]\\ &=\mathbb{E}[X]-\frac{\operatorname{Cov}\left(X,\Psi^{-1}(1-F_{\mu}(X))\right)}{\operatorname{Var}(h(Z))}\mathbb{E}[h(Z)],\end{split}$
Now we can prove Theorem 1. By the linearity of covariances,
$\begin{split}&\operatorname{Cov}\left(X,\Psi^{-1}(F_{\mu}(X))\right)+\operatorname{Cov}\left(X,\Psi^{-1}(1-F_{\mu}(X))\right)\\ =&\operatorname{Cov}\left(X,\Psi^{-1}(F_{\mu}(X))+\Psi^{-1}(1-F_{\mu}(X))\right).\end{split}$
If (8) holds, the above implies that the value in (43) is smaller or equal to that in (46). Hence, the minimizer $(\theta_{1}^{*},\theta_{2}^{*})$ should be the one associated with the value (43), which is given by (9). On the other hand, if (8) fails, the same argument shows that the value in (46) is smaller than that in (43). The minimizer $(\theta_{1}^{*},\theta_{2}^{*})$ should then be the one associated with the value (46), which is given by (45).
@@ IV-B Proof for Theorem 3
In the beginning, here we provide the proof sketch of Lemma 1 and left the details to Appendix C. First, we assume $X$ is zero mean, since equivalently one can modify $G_{\Theta}(Z)$ in (II) with a bias $G_{\Theta}(Z)+\mathbb{E}[X]$ for non-zero mean $X$. With $q=2$, Wasserstein distance (as RHS of (2)) of order 2 between $\omega^{T}X$ with $\hat{X}_{G}$, averaged over $\omega$, is bounded from [17]. We modify this result to make it valid for $q=1$, and then get desired bounded gap $f^{\mu}(d)$ between (21)(22) for $q=1,2$ as
@@
where $C^{\prime}$ is a constant. With linear generator (3), not only $\omega^{T}\Theta Z$ is Gaussian as $\hat{X}_{G}$ given $\omega$, on the contrary to [16, Theorem 1], but also we prevent gap $f^{\mu}(d)$ depending on $\Theta$.
@@
where $M_{U_{w}}(z):=\mathbb{E}[e^{zU_{w}}]$ is the moment generating function of $U_{w}$. For the second term it is easy to see that
@@
Combining these results reach (24).
Continuing the proof for Theorem 3, we first aim on solving (24). Though the derivative with respect to $z$ in (24) can be easily solved by matrix calculus [26], directly integration over $z$ is still too complicated for turn it to the closed-form (27). To solve this, applying SVD on the parameters of linear generator $\Theta=\mathbf{U}\Lambda\mathbf{V}^{T}\in\mathbb{R}^{d\times r}$ then $\Theta\Theta^{T}$ can be eigen-decomposed as
@@
$=\frac{\sum_{i=1}^{d}S_{ii}}{d}-\frac{2\tilde{\sigma}_{x}}{\Gamma(1/2)}\int^{\infty}_{0}\frac{1}{2}z^{-1/2}(\prod_{i=1}^{d}(1+\frac{2z}{d}S_{ii})^{-3/2})$
$(\sum_{i=1}^{d}\frac{2}{d}S_{ii}\prod_{j=1,j\neq i}^{d}(1+\frac{2z}{d}S_{jj}))dz.$ (53)
To solve the optimal elements of the diagonal matrix $\mathbf{S}$ in (53), we focus on simplifying the integral part of it using the Carlson-R function and turn it to objectives of (27). Details are as follows, in (53),
$-\frac{2\tilde{\sigma}_{x}}{\Gamma(1/2)}\int^{\infty}_{0}\frac{1}{2}z^{-1/2}(\prod_{i=1}^{d}(1+\frac{2z}{d}S_{ii})^{-3/2})$
$(\sum_{i=1}^{d}\frac{2}{d}S_{ii}\prod_{j=1,j\neq i}^{d}(1+\frac{2z}{d}S_{jj}))dz$
$=$ $-\frac{2\tilde{\sigma}_{x}}{\Gamma(1/2)}\sum_{i=1}^{d}\int^{\infty}_{0}\frac{1}{d}z^{-1/2}S_{ii}(1+\frac{2z}{d}S_{ii})^{-3/2}$
$\prod_{j=1,j\neq i}^{d}(1+\frac{2z}{d}S_{jj})^{-1/2}dz.$
@@
The step (a) follows [18, (8.3-1)] where $B(\cdot,\cdot)$ is beta function, $\mathbf{b}^{(i)}\in\mathbb{R}^{d}$ with elements $b_{i}^{(i)}=3/2$ and $b_{j}^{(i)}=1/2$ for $j\neq i$, and $R_{-a}(\mathbf{b};\mathbf{z})$ is the Carlson-R function
$R_{-a}(\mathbf{b};\mathbf{z})$ $=\frac{1}{B(a,\sum{b_{i}})}\int_{0}^{\infty}t^{\sum{b_{i}}-a}\prod(t+z_{i})^{-b_{i}}dt$
$=\frac{1}{B(a,\sum{b_{i}})}\int_{0}^{\infty}t^{a-1}\prod(1+tz_{i})^{-b_{i}}dt.$
@@
the step (b) follows [18, (5.9-3)] that
@@
and the step (c) comes from Euler’s homogeneity relation [18, (5.9-2)]
@@
Next, with aids of Dirichlet average [18], we prove that the objective of (27) is Schur-convex [19], and it achieves the global minimum when all the diagnol elements of $\mathbf{S}$ are the same as $s$.
@@
The objective of equivalent problem (27) of (24) is symmetric and convex over $\mathbf{S}$, with such a Schur-convex objective (27) has global minimum when $\mathbf{S}=s\mathbf{I}$.
@@
Carlson R function is a special case of Dirichlet average[18, (5.9-1)]
@@
Note that (55) is symmetric over $\mathbf{S}$ obviously, and thus same property applies to (27). Since the Hessian matrix of $(\mathbf{u}\cdot\mathrm{diag}(\mathbf{S}))^{1/2}$ is positive semidefinite and the nonnegative weighted integral preserves the convexity, (55) is concave and thus (27) is convex over $\mathbf{S}$. ∎
@@
$\min_{s}s-\tilde{\sigma}_{x}\sqrt{2d}\frac{\Gamma({\frac{d+1}{2})}}{\Gamma({\frac{d}{2}+1)}}R_{\frac{1}{2}}(\frac{1}{2}\mathbf{1};s\cdot\mathbf{1})$ (56)
$=$ $\min_{s}s-\tilde{\sigma}_{x}\sqrt{2d}\frac{\Gamma({\frac{d+1}{2})}}{\Gamma({\frac{d}{2}+1)}}s^{1/2},$
where the equality follows [27, 19.16.11]. Then, by solving the quadratic programming the optimal
@@
By Lemma 1, the upper bound (26) for (II) from (III-B) is obtained. From [27, (5.11.12)], we have the asymptotic equivalence (same notation as $X\sim\mathcal{N}$)
@@
The upper bound (26) then goes to zero as $d\rightarrow\infty$, and the lower bound of (II) is zero obviously. Thus, the linear generator with the parameters (19) is asymptotically optimal for the sliced WGAN (II).

### v2

@@ IV-A Proof for Theorem 1
We first show that the covariance term in (37) is non-negative. As $\Phi^{-1}$ is strictly increasing and $F_{\mu}$ is nondecreasing, the map $x\mapsto\Phi^{-1}(F_{\mu}(x))1_{\{F_{\mu}(x)>1/2\}}$ is nondecreasing. This readily implies
$\operatorname{Cov}\left(X,\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)>1/2\}}\right)\geq 0$ (49)
in (37) since
$(x-x^{\prime})\big(\Phi^{-1}(F_{\mu}(x))1_{\{F_{\mu}(x)>1/2\}}-\Phi^{-1}(F_{\mu}(x^{\prime}))1_{\{F_{\mu}(x^{\prime})>1/2\}}\big)\geq 0$
for all $x,x^{\prime}\in\mathbb{R}$. Specifically, by taking an independent but same distribution copy $X^{\prime}$ of $X$,
$0$ $\leq\mathbb{E}[(X-X^{\prime})\cdot\left(\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)>1/2\}}-\Phi^{-1}(F_{\mu}(X^{\prime}))1_{\{F_{\mu}(X^{\prime})>1/2\}}\right)]$ (50)
$=2\operatorname{Cov}\left(X,\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)>1/2\}}\right).$
Specifically, we deal with the two cases where $\operatorname{Cov}\left(X,\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)>1/2\}}\right)$ is strictly positive and zero separately.
Case I: $\operatorname{Cov}\left(X,\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)>1/2\}}\right)>0$. If $\theta_{2}=0$, (37) entails
$\lambda=-2\operatorname{Cov}\left(X,\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)>1/2\}}\right)<0$, which violates the requirement $\lambda\geq 0$ in (38)-(39). If $\lambda=0$, solving (37) yields (48).
That is, the KKT condition gives a unique candidate optimizer $(\theta_{1}^{*},\theta_{2}^{*},\lambda^{*})$, with $(\theta_{1}^{*},\theta_{2}^{*})$ as in (48) and $\lambda^{*}=0$. To check the corresponding second-order condition, let $H_{J}$ and $H_{g}$ denote the Hessian matrices of $J(\theta_{1},\theta_{2})$ and $g(\theta_{1},\theta_{2})$, respectively. For $H_{J}$,
$\frac{\partial^{2}J(\theta_{1},\theta_{2})}{{\partial\theta_{1}}^{2}}=2,$
$\frac{\partial^{2}J(\theta_{1},\theta_{2})}{\partial\theta_{1}\partial\theta_{2}}=2\mathbb{E}\left[\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)>1/2\}}\right],$
from (46) and
$\frac{\partial^{2}J(\theta_{1},\theta_{2})}{{\partial\theta_{2}}^{2}}=2\mathbb{E}\left[(\Phi^{-1}(F_{\mu}(X)))^{2}1_{\{F_{\mu}(X)>1/2\}}\right],$
from (47). Clearly, $H_{g}=O_{2\times 2}$, which implies
$H_{J}+\lambda H_{g}=2\begin{bmatrix}1&\frac{1}{\sqrt{2\pi}}\\ \frac{1}{\sqrt{2\pi}}&\frac{1}{2}\end{bmatrix}$, (51)
which is positive definite. Hence, we conclude that $(\theta_{1}^{*},\theta_{2}^{*})$ in (48) is the unique minimizer of (36).
Case II: $\operatorname{Cov}\left(X,\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)>1/2\}}\right)=0$. As (37) entails $\lambda=\theta_{2}(1-\frac{1}{\pi})$, we have $\lambda=\theta_{2}=0$ (by recalling complementary slackness that either $\lambda=0$ or $\theta_{2}=0$). That is, the KKT condition gives a unique candidate optimizer $(\theta_{1}^{*},\theta_{2}^{*},\lambda^{*})=(\mathbb{E}[X],0,0)$. Since $H_{J}+\lambda H_{g}$ is positive definite, we conclude that $(\theta_{1}^{*},\theta_{2}^{*})=(\mathbb{E}[X],0)$ is the unique minimizer of (36).
As a remark, another way to prove is to show $H_{J}$ as (51) is positive definite in both cases first, and then $J(\theta_{1},\theta_{2})$ is convex. Since both proofs require checking the Hessian matrix, there is no simpler way to prove. ∎
Lemma 5.
$\begin{split}&\int_{\mathbb{R}}|x-t^{\Theta^{*}}(x)|^{2}d\mu(x)=\operatorname{Var}(X)-\frac{2\pi}{\pi-1}\operatorname{Cov}\left(X,\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)>1/2\}}\right)^{2}.\end{split}$ (52)
Proof.
@@
$=$ $\int_{\{F_{\mu}(x)\leq 1/2\}}\left(x-\mathbb{E}[X]+\frac{\theta^{*}_{2}}{\sqrt{2\pi}}\right)^{2}d\mu(x)+\int_{\{F_{\mu}(x)>1/2\}}\left(x-\mathbb{E}[X]-\frac{\theta^{*}_{2}}{\sqrt{2\pi}}-\theta^{*}_{2}\Psi^{-1}(F_{\mu}(x))\right)^{2}d\mu(x)$
$=$ $\int_{\{F_{\mu}(x)\leq 1/2\}}(x-\mathbb{E}[X])^{2}d\mu(x)+\sqrt{\frac{2}{\pi}}\theta^{*}_{2}\int_{\{F_{\mu}(x)\leq 1/2\}}(x-\mathbb{E}[X])d\mu(x)+\frac{(\theta^{*}_{2})^{2}}{2\pi}\int_{\mathbb{R}}1_{\{F_{\mu}(x)\leq 1/2\}}d\mu(x)$
$+$ $\int_{\{F_{\mu}(x)>1/2\}}(x-\mathbb{E}[X])^{2}d\mu(x)-2\theta^{*}_{2}\int_{\{F_{\mu}(x)>1/2\}}(x-\mathbb{E}[X])\Big(\Psi^{-1}(F_{\mu}(x))-\frac{1}{\sqrt{2\pi}}\Big)d\mu(x)$
Observe that
$=$ $\int_{\mathbb{R}}(x-\mathbb{E}[X])\Big(\Psi^{-1}(F_{\mu}(x))1_{\{F_{\mu}(x)>1/2\}}-\frac{1}{\sqrt{2\pi}}1_{\{F_{\mu}(x)>1/2\}}\Big)d\mu(x)$
$=$ $\int_{\mathbb{R}}(x-\mathbb{E}[X])\Big(\Psi^{-1}(F_{\mu}(x))1_{\{F_{\mu}(x)>1/2\}}-\frac{1}{\sqrt{2\pi}}\Big)d\mu(x)+\frac{1}{\sqrt{2\pi}}\int_{\mathbb{R}}(x-\mathbb{E}[X])1_{\{F_{\mu}(x)\leq 1/2\}}d\mu(x)$
$=$ $\operatorname{Cov}\left(X,\Psi^{-1}(F_{\mu}(x))1_{\{F_{\mu}(x)>1/2\}}\right)+\frac{1}{\sqrt{2\pi}}\int_{\mathbb{R}}(x-\mathbb{E}[X])1_{\{F_{\mu}(x)\leq 1/2\}}d\mu(x),$ (54)
where the last equality follows from (46). Similarly,
$=$ $\int_{\mathbb{R}}\Big(\Psi^{-1}(F_{\mu}(x))1_{\{F_{\mu}(x)>1/2\}}-\frac{1}{\sqrt{2\pi}}\Big)^{2}d\mu(x)-\frac{1}{2\pi}\int_{\mathbb{R}}1_{\{F_{\mu}(x)\leq 1/2\}}d\mu(x)$
$\int_{\mathbb{R}}|x-t^{\Theta^{*}}(x)|^{2}d\mu(x)$
$=$ $\operatorname{Var}(X)-2\theta^{*}_{2}\operatorname{Cov}\left(X,\Psi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(x)>1/2\}}\right)+(\theta^{*}_{2})^{2}\frac{\pi-1}{2\pi}$
where the last equality is due to $\theta^{*}_{2}$ in (48). ∎
Next, we go to $\theta_{2}\leq 0$. The target sub-problem is
$\min_{\theta_{1}\in\mathbb{R},\theta_{2}\leq 0}\int_{\mathbb{R}}|x-t^{\Theta}(x)|^{2}d\mu(x).$ (56)
@@
$\int_{\mathbb{R}}|x-t^{\Theta^{*}}(x)|^{2}d\mu(x)=\operatorname{Var}(X)-\frac{2\pi}{\pi-1}\operatorname{Cov}\left(X,\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)\leq 1/2\}}\right)^{2}.$ (58)
The proof for the case $\theta_{2}\leq 0$ is similar to the case $\theta_{2}\geq 0$, and is given in Appendix B Finally, we can get our theorem. If (13) holds, the value in (52) is smaller than or equal to that in (58). Hence, the minimizer $(\theta_{1}^{*},\theta_{2}^{*})$ should be the one associated with the value (52), which is given by (48). On the other hand, if (13) fails, the value in (58) is smaller than that in (52). The minimizer $(\theta_{1}^{*},\theta_{2}^{*})$ should then be the one associated with the value (58), which is given by (57).
@@ IV-B Proof for Theorem 2
In the beginning, here we provide the proof sketch of Lemma 1 and leave the details to Appendix D. First, we assume $X$ is zero mean, since equivalently one can modify $G_{\Theta}(Z)$ in (6) with a bias $G_{\Theta}(Z)+\mathbb{E}[X]$ for non-zero mean $X$. With $q=2$, Wasserstein distance (as RHS of (2)) of order 2 between $\omega^{T}X$ with $\hat{X}_{G}$, averaged over $\omega$, is bounded from [28]. We modify this result to make it valid for $q=1$, and then get desired bounded gap $f^{\mu}(d)$ between (22)(23) for $q=1,2$ as
@@
where $C^{\prime}$ is a constant. With linear generator (3), not only $\omega^{T}\Theta Z$ is Gaussian as $\hat{X}_{G}$ given $\omega$, on the contrary to [25, Theorem 1], but also we prevent gap $f^{\mu}(d)$ depending on $\Theta$.
@@
where $M_{U_{w}}(z):=\mathbb{E}[e^{zU_{w}}]$ is the moment generating function of $U_{w}$. For the second term, it is easy to see that
@@
Combining these results reaches (25).
Continuing the proof for Theorem 2, we first aim to solve (25). Though the derivative with respect to $z$ in (25) can be easily solved by matrix calculus [15], direct integration over $z$ is still too complicated to turn it into the closed-form (28). To solve this, applying SVD on the parameters of linear generator $\Theta=\mathbf{U}\Lambda\mathbf{V}^{T}\in\mathbb{R}^{d\times r}$, then $\Theta\Theta^{T}$ can be eigen-decomposed as
@@
$=$ $\frac{\sum_{i=1}^{d}S_{ii}}{d}-\frac{2\tilde{\sigma}_{x}}{\Gamma(1/2)}\int^{\infty}_{0}\frac{1}{2}z^{-1/2}(\prod_{i=1}^{d}(1+\frac{2z}{d}S_{ii})^{-3/2})(\sum_{i=1}^{d}\frac{2}{d}S_{ii}\prod_{j=1,j\neq i}^{d}(1+\frac{2z}{d}S_{jj}))dz.$ (64)
To solve the optimal elements of the diagonal matrix $\mathbf{S}$ in (64), we focus on simplifying the integral part of it using the Carlson-R function and turn it into the objectives of (28). Details are as follows, in (64),
$-\frac{2\tilde{\sigma}_{x}}{\Gamma(1/2)}\int^{\infty}_{0}\frac{1}{2}z^{-1/2}(\prod_{i=1}^{d}(1+\frac{2z}{d}S_{ii})^{-3/2})(\sum_{i=1}^{d}\frac{2}{d}S_{ii}\prod_{j=1,j\neq i}^{d}(1+\frac{2z}{d}S_{jj}))dz$ (65)
$=$ $-\frac{2\tilde{\sigma}_{x}}{\Gamma(1/2)}\sum_{i=1}^{d}\int^{\infty}_{0}\frac{1}{d}z^{-1/2}S_{ii}(1+\frac{2z}{d}S_{ii})^{-3/2}\prod_{j=1,j\neq i}^{d}(1+\frac{2z}{d}S_{jj})^{-1/2}dz.$
@@
The step (a) follows [7, (8.3-1)] where $B(\cdot,\cdot)$ is beta function, $\mathbf{b}^{(i)}\in\mathbb{R}^{d}$ with elements $b_{i}^{(i)}=3/2$ and $b_{j}^{(i)}=1/2$ for $j\neq i$, and $R_{-a}(\mathbf{b};\mathbf{z})$ is the Carlson-R function
$R_{-a}(\mathbf{b};\mathbf{z})$ $=\frac{1}{B(a,\sum{b_{i}})}\int_{0}^{\infty}t^{\sum{b_{i}}-a}\prod(t+z_{i})^{-b_{i}}dt=\frac{1}{B(a,\sum{b_{i}})}\int_{0}^{\infty}t^{a-1}\prod(1+tz_{i})^{-b_{i}}dt.$
@@
the step (b) follows [7, (5.9-3)] that
@@
and the step (c) comes from Euler’s homogeneity relation [7, (5.9-2)]
@@
Next, with the aid of Dirichlet average [7], we prove that the objective of (28) is Schur-convex [22], and it achieves the global minimum when all the diagonal elements of $\mathbf{S}$ are the same as $s$.
@@
The objective of the equivalent problem (28) of (25) is symmetric and convex over $\mathbf{S}$, and such a Schur-convex objective (28) has a global minimum when $\mathbf{S}=s\mathbf{I}$.
@@
Carlson R function is a special case of Dirichlet average[7, (5.9-1)]
@@
Note that (67) is symmetric over $\mathbf{S}$ obviously, and thus the same property applies to (28). Since the Hessian matrix of $(\mathbf{u}\cdot\mathrm{diag}(\mathbf{S}))^{1/2}$ is positive semidefinite and the nonnegative weighted integral preserves the convexity, (67) is concave and thus (28) is convex over $\mathbf{S}$. ∎
@@
$\min_{s}s-\tilde{\sigma}_{x}\sqrt{2d}\frac{\Gamma({\frac{d+1}{2})}}{\Gamma({\frac{d}{2}+1)}}R_{\frac{1}{2}}(\frac{1}{2}\mathbf{1};s\cdot\mathbf{1})=\min_{s}s-\tilde{\sigma}_{x}\sqrt{2d}\frac{\Gamma({\frac{d+1}{2})}}{\Gamma({\frac{d}{2}+1)}}s^{1/2},$ (68)
where the equality follows [26, 19.16.11]. Then, by solving the quadratic programming, the optimal
@@
By Lemma 1, the upper bound (27) for (6) from (III-B) is obtained. From [26, (5.11.12)], we have the asymptotic equivalence (same notation as $X\sim\mathcal{N}$)
@@
The upper bound (27) then goes to zero as $d\rightarrow\infty$, and the lower bound of (6) is obviously zero. Thus, the linear generator with the parameters (20) is asymptotically optimal for the sliced WGAN (6).